### v1

# Non-Convex Stochastic Composite Optimization with Polyak Momentum

Abstract
The stochastic proximal gradient method is a powerful generalization of the widely used stochastic gradient descent (SGD) method and has found numerous applications in Machine Learning. However, it is notoriously known that this method fails to converge in non-convex settings where the stochastic noise is significant (i.e. when only small or bounded batch sizes are used). In this paper, we focus on the stochastic proximal gradient method with Polyak momentum. We prove this method attains an optimal convergence rate for non-convex composite optimization problems, regardless of batch size. Additionally, we rigorously analyze the variance reduction effect of the Polyak momentum in the composite optimization setting and we show the method also converges when the proximal step can only be solved inexactly. Finally, we provide numerical experiments to validate our theoretical results.
1 Introduction
Stochastic gradient descent and its variants are the workhorse of modern machine learning. The stochastic proximal gradient method is a simple yet powerful extension of the vanilla stochastic gradient descent method, which aims to solve the following stochastic composite optimization problem:
$\min_{\mathbf{x}\in\mathbb{R}^{d}}\{F(\mathbf{x})\coloneqq f(\mathbf{x})+\psi(\mathbf{x})\}\,,$ (1)
where $f:\mathbb{R}^{d}\to\mathbb{R}$ is smooth and $\psi\colon\mathbb{R}^{d}\to\mathbb{R}\cup\{+\infty\}$ is not necessarily differentiable, but a simple function. Such paradigm is ubiquitous in machine learning and beyond, and it covers a wide range of generalizations of the vanilla optimization problem, including regularized machine learning problems [26], signal processing [6], image processing [28] and many more. It also naturally covers constrained optimization problems by considering the indicator function of the constraint set, and some recent variants of such ideas have been applied to distributed and federated machine learning problems [29].
The problem formulation (1) is often used in Machine Learning as the proximal term $\phi$ allows encoding prior domain knowledge. For instance, it can impose constraints on the variables, handle non-differentiability, induce sparsity, or it can take the form of a regularizer. Recently, there has been a surge of interest in these preconditioning techniques in the Deep Learning context [17, 41]. Given a loss objective function $\ell\colon\mathbb{R}^{d}\to\mathbb{R}$, an auxiliary function (or preconditioner) $\hat{\ell}$ is constructed—for instance, by approximating $\ell$ with a subset of the data samples or synthetic samples. This results in the formulation:
$\min_{\mathbf{x}\in\mathbb{R}^{d}}\{\underbrace{\ell(\mathbf{x})-\hat{\ell}(\mathbf{x})}_{f(\mathbf{x})}+\underbrace{\hat{\ell}(\mathbf{x})}_{\psi(\mathbf{x})}\}\,.$
This is a special case of Problem (1), with $f\coloneqq\ell-\hat{\ell}$ and $\psi\coloneqq\hat{\ell}$. This formulation has the advantage that the training can be accelerated if the proxy function $\hat{\ell}$ remains simple to optimize. These applications motivate the need to further our understanding of the stochastic composite optimization problem in the non-convex regime, especially where large batches are less preferred or even unavailable [24, 34].
In the convex and strongly convex case, the complexity of solving Problem (1) is well understood [16, 14]. On the other hand, the existing theory in the non-convex regime is unsatisfactory. In particular, when the gradient noise has variance $\sigma^{2}$, with the vanilla algorithm, the squared norm of the gradient of $F$ is only shown to converge to $\mathcal{O}(\nicefrac{{1}}{{K}})+\Omega(\sigma^{2})$ after $K$ iterations. This implies that to converge to an arbitrary $\varepsilon$ error, one needs to take mega batches of size $\Omega(\varepsilon^{-1})$ at each iteration of the algorithm, and the total number of stochastic gradient oracle calls is $\Omega(\varepsilon^{-2})$. In practice and theory, smaller-batch methods are often preferred over mega-batch methods [^1] By a mega-batch we mean a batch of the size of the order $\Omega(\varepsilon^{-1})$.. For example, mega-batch methods follow the full gradient methods much more closely at each step than smaller batch methods, and it is observed empirically to be adversarial to the generalization performance [40, 24, 20], and theoretically [36] for certain Machine Learning tasks. Furthermore, in many practical settings, mega-batches are unavailable or intractable to sample, e.g., in medical tasks [34]; federated Reinforcement Learning [21, 19]; and multi-agent Reinforcement Learning [10].
In this work, we revisit the stochastic composite optimization problem in the non-convex regime and show that the Polyak momentum technique addresses the aforementioned mega-batch issue while retaining the optimal convergence rate $\mathcal{O}(\varepsilon^{-2})$. Polyak momentum is the de-facto standard for training Deep Learning models in practice [22], and understanding its theoretical efficacy in various settings is an active research direction [7, 12].
1.1 Stochastic proximal gradient method
To solve the stochastic composite optimization problem (1), we consider the stochastic proximal gradient method, which solves the following subproblem in each iteration (commonly known as the proximal step):
$\mathbf{x}_{k+1}$ $\coloneqq\operatorname*{arg\,min}_{\mathbf{x}\in\mathbb{R}^{d}}\{\Omega_{k}(\mathbf{x})\}\,,\text{with}$ (2)
$\Omega_{k}(\mathbf{x})$ $\coloneqq\langle\mathbf{g}_{k},\mathbf{x}\rangle+\psi(\mathbf{x})+\frac{M}{2}\left\lVert\mathbf{x}-\mathbf{x}_{k}\right\rVert^{2}\,.$
Here, $\mathbf{g}_{k}$ denotes a random vector computed at each iteration $k$ using the stochastic first-order oracle of $f$, and $M>0$ is a regularization (or stepsize) parameter. For the vanilla stochastic proximal gradient method, $\mathbf{g}_{k}$ is a stochastic gradient of $f$ at $\mathbf{x}_{k}$. We note that when $\psi\equiv 0$, this reduces to the vanilla SGD with step size $\frac{1}{M}$. When $\psi\not\equiv 0$, it has long been known that if error-dependent batches are not allowed, then such a method is only proved to converge to a neighborhood of the stationary point of $F$ up to the variance of the gradient noise [16]. In Section 4 we give a simple example to illustrate this phenomenon. Therefore, one needs $\varepsilon$-dependent mini-batches (or mega-batches) to converge to an $\varepsilon$ neighborhood of the stationary point of $F$.
1.2 Our contributions
Stochastic Polyak momentum has seen a huge success in practice [22]. In this work, we study the effect of Polyak momentum in the non-convex regime and provide a theoretical analysis of momentum’s stabilizing effect in the stochastic composite optimization setting, particularly in the small batch regime.
  • First, we establish a lower bound result for the vanilla stochastic proximal gradient method, showing that it cannot converge to the stationary point beyond the variance of the gradient noise.
  • We study the effect of incorporating momentum into the stochastic proximal gradient method and prove its optimal convergence to the stationary point without any error-dependent mega-batch access. We also rigorously study the variance reduction effect of momentum in the stochastic composite optimization setting.
  • We further extend our analysis to the case where the proximal steps are solved inexactly and give the same convergence guarantees, demonstrating the robustness of the momentum method.
  • Finally, we conduct numerical experiments to corroborate our theoretical findings and demonstrate the algorithm’s practicality.
2 Related works
There is a huge body of work on stochastic and composite optimization. Here, we focus on the stochastic and non-convex regime and do not get into the details of the works in convex (e.g. [13, 15]) or deterministic regime (see [31]). The first work that considers the stochastic composite optimization problem in the non-convex case appears to be [16], in which the convergence of the stochastic proximal gradient method to a neighborhood of the stationary point of $F$ up to the variance of the gradient noise is established. If mega-batches are allowed, the paper shows that the algorithm takes asymptotically $\mathcal{O}(\varepsilon^{-2})$ stochastic gradient oracle calls to converge to an $\varepsilon$ neighborhood of the stationary point of $F$. These results were later extended to incorporate the acceleration technique, but the convergence still requires mega-batches. In particular, the method requires $k$ samples at the $k$th iteration while the total number of oracle calls is not improved [14]. Even though these methods require mega-batches, their upper bounds on the total number of stochastic gradient oracle calls match the lower bound asympotically [2].
Average smoothness and variance reduction
In another research direction, to break the $\mathcal{O}(\varepsilon^{-2})$ lower bound for stochastic non-convex optimization, many variance reduction techniques have been proposed, at the cost of two additional assumptions on the objective and problem structure:
  • $f$ is $\widetilde{L}$-average smooth: ${\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x},\xi)-\nabla f(\mathbf{y},\xi)\right\rVert^{2}\right]\leq\widetilde{L}^{2}\left\lVert\mathbf{x}-\mathbf{y}\right\rVert^{2}$. Note that this assumption might not be satisfied for some simple smooth functions, and it is strictly stronger than the standard smoothness assumption. [^2] By Jensen’s inequality, we have that any $\widetilde{L}$-average smooth function is also $\widetilde{L}$-smooth, see also 3.3. There exists $L$-smooth functions that are not average smooth.
  • These variance reduction techniques require the access to two stochastic gradients $\nabla f(\mathbf{x}_{k},\xi_{k})$ and $\nabla F(\mathbf{x}_{k-1},\xi_{k})$ at iteration $k$, which might not always be available in practice (e.g. [10, 5]).
There have been a long line of works on variance reduction techniques, including Prox-Spiderboost [39] (composite variant of Spider [11]), Hybrid-SGD [37], and PStorm [42] (composite variant of Storm [8]). These methods achieve a $\mathcal{O}(\varepsilon^{-\nicefrac{{3}}{{2}}})$ asymptotic complexity.
Convergence criteria
The first-order convergence criteria for non-convex optimization problems is an ongoing research topic. In most of the works studying the non-convex composite optimization problem (1), the convergence is studied in terms of the proximal gradient mapping [16, 14, 39, 37, 42]. There have been several works proposing different convergence criteria for non-convex optimization problems and discussing how these definitions affect the corresponding convergence complexities, e.g. [9, 44, 45], but these are parallel to our work, and we do not discuss them in detail here. Our work mostly focuses on the convergence in terms of $\left\lVert\nabla F(\mathbf{x}_{k})\right\rVert^{2}$. We note that all out theory works with proximal gradient mapping criteria, see Appendix D.
Stochastic polyak momentum
The idea of using Polyak momentum (a.k.a. heavy-ball momentum) was first proposed in [32] for strongly convex quadratic objective in the deterministic case. The first non-asymptotic analysis of SGD with Polyak momentum in the smooth non-convex regime is given in [43] and was later refined in [27]. Some recent works also studied how Polyak momentum can be used to remove the dependence on large batches in various settings, e.g. for normalized SGD [7] and for communication compressed optimization [12]. The analysis of stochastic gradient methods with Polyak momentum remains an active research topic [38, 35, 25, 18].
3 Problem formulation and assumptions
In this work, we consider the composite optimization problem:
$\min_{\mathbf{x}\in\mathbb{R}^{d}}\{F(\mathbf{x})\coloneqq f(\mathbf{x})+\psi(\mathbf{x})\}\,.$
We study the convergence of the algorithms in terms of $\left\lVert\nabla F(\mathbf{x})\right\rVert^{2}$. When $\psi$ is convex and non-differentiable, $\nabla F$ is understood as an appropriate choice of the subgradient. We do not consider the non-convex and non-differentiable case for $\psi$, as this is an exotic scenario in the literature. For simplicity, we assume differentiability when presenting the proofs in the main text, but all our results hold for the convex and non-differentiable case for $\psi$.
Note that most of the existing works on stochastic composite optimization assume that $\psi$ is convex and analyze the convergence of their algorithms in terms of the proximal gradient mapping [16, 14, 39, 37, 42], which is closely related to what we consider in this work. In Appendix D, we discuss the proximal gradient mapping and argue that our definition is more natural in the non-convex regime.
We first introduce the following assumption on the monotonicity of the proximal step:
Assumption 3.1.
We assume that, at each iteration $k$ of the algorithms, the output $\mathbf{x}_{k+1}$ of the proximal step (2) satisfies the following:
$\Omega_{k}(\mathbf{x}_{k+1})\leq\Omega_{k}(\mathbf{x}_{k})\,.$
This is a technical assumption that provides a more nuanced characterization of $\psi$ beyond convexity, but all our theories work with just the simple convexity assumption as well (in fact, several constants can be improved under the convexity assumption). It is natural to assume that the proximal step finds a no worse point than the current iterate, even when $\psi$ is non-convex.
Next, we introduce the lower boundedness assumption of the objective function $F$:
Assumption 3.2.
We assume that there exists $F^{\star}\in\mathbb{R}$ such that $F^{\star}\leq F(\mathbf{x}),\forall\mathbf{x}\in\mathbb{R}^{d}$.
Now, we introduce the smoothness assumption on $f$. Note that this is a weaker assumption than the average smoothness assumption used in the variance reduction type method [39, 37, 42].
Assumption 3.3.
We assume that the function $f$ has $L$-Lipschitz gradient, i.e. for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, we have:
$\left\lVert\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\rVert\leq L\left\lVert\mathbf{x}-\mathbf{y}\right\rVert\,.$
Next, we make an assumption on the noise of the gradient oracle of $f$.
Assumption 3.4.
We assume that we have access to a gradient oracle $\nabla f(\mathbf{x},\xi)$ for $f$ such that for all $\mathbf{x}\in\mathbb{R}^{d}$ it holds that:
${\mathbb{E}}\left[\nabla f(\mathbf{x},\xi)\right]=\nabla f(\mathbf{x})\,,$ (3)
${\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x},\xi)-\nabla f(\mathbf{x})\right\rVert^{2}\right]\leq\sigma^{2}.$
This is a standard assumption in the analysis of stochastic gradient methods [30, 4]. Taking mini-batches is equivalent to dividing the variance by the batch size.
4 Lower bound for the vanilla stochastic proximal gradient method
In Section 1.1 we briefly discussed that the vanilla stochastic proximal gradient method cannot converge to the stationary point of $F$ beyond the variance of the gradient noise. We give a simple example to illustrate this phenomenon and give some intuitions on why momentum resolves the issue.
Proposition 4.1.
For any step-size coefficient $M>0$, there exists $F=f+\psi:\mathbb{R}^{d}\to\mathbb{R}$ with $L$-smooth $f$ and stochastic gradient oracle of $f$ with variance $\sigma^{2}$, such that the method (2) with $\mathbf{g}_{k}$ being the stochastic gradient of $f$ at $\mathbf{x}_{k}$ cannot converge to the stationary point of $F$ beyond a $\frac{\sigma^{2}}{4}$ neighborhood. In particular, given any $\mathbf{x}_{k}$, ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\geq\frac{\sigma^{2}}{4}$.
Proof.
Consider $f(\mathbf{x})=\frac{L\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ and $\psi(\mathbf{x})=\frac{a\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ for some $a>0$. The stochastic gradient oracle of $f$ is $\mathbf{g}_{k}=\nabla f(\mathbf{x}_{k})+\xi_{k}$ where $\xi_{k}\sim\mathcal{N}(0,\sigma^{2}\mathbf{I})$. Clearly, the construction satisfies 3.3 and 3.4. It is easy to check that the proximal step is given by:
$\mathbf{x}_{k+1}=\frac{M-L}{M+a}\mathbf{x}_{k}-\frac{\xi_{k}}{M+a}\,.$
Therefore,
${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $=\frac{(M+L)^{2}}{(M+a)^{2}}{\mathbb{E}}\left[\left\lVert(M-L)\mathbf{x}_{k}-\xi_{k}\right\rVert^{2}\right]$
$=\frac{(M+L)^{2}}{(M+a)^{2}}\left((M-L)^{2}\left\lVert\mathbf{x}_{k}\right\rVert^{2}+\sigma^{2}\right)$
$\geq\frac{(M+L)^{2}\sigma^{2}}{(M+a)^{2}}\,,$
where the first equality follows from the independence of $\xi_{i}$ and $\mathbf{x}_{k}$. Hence, for any $a>M$, we must have that method (2) with the stochastic gradient $\mathbf{g}_{k}$ cannot converge to the stationary point of $F$ beyond $\frac{\sigma^{2}}{4}$ neighborhood. ∎
Remark 4.2.
We point out that one of the characteristics of the proximal gradient method is that it is oblivious to the choice of $\psi$. In particular, from a theoretical point of view, the algorithm is only aware of the parameters of $f$, and the convergence guarantees should always hold despite of the choice of $\psi$.
The problem is that the variance of the gradient noise keeps the iterates away from the stationary point of $F$, and we need some mechanism to reduce the variance of the gradient noise towards the end of the optimization process. This is when the Polyak momentum comes into play.
5 The algorithm and analysis
In this section, we present the stochastic proximal gradient method with Polyak momentum and give its convergence guarantees in the non-convex regime.
Algorithm 1 Proximal Gradient Method with Polyak Momentum
[…]
6:$\Omega_{k}(\mathbf{x})\coloneqq\langle\mathbf{m}_{k},\mathbf{x}\rangle+\psi(\mathbf{x})+\frac{M}{2}\left\lVert\mathbf{x}-\mathbf{x}_{k}\right\rVert^{2}$
7:s.t. $\Omega_{k}(\mathbf{x}_{k+1})\leq\Omega_{k}(\mathbf{x}_{k})$ and $\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\leq\delta_{k}$
[…]
The algorithm is given in Algorithm 1. In the algorithm, the stochastic gradient $\mathbf{g}_{k}$ at each iteration $k$ is replaced by the momentum - aggregated gradient mean:
$\mathbf{m}_{k}=(1-\gamma)\mathbf{m}_{k-1}+\gamma\mathbf{g}_{k}\,.$
Our analysis will show that the distance between the momentum and the full gradient decreases as the number of iterations increases, which is the key to resolving the issue of the vanilla stochastic proximal gradient method.
In this section, we assume that the proximal steps are solved exactly, i.e. we further make the following assumption:
Assumption 5.1.
At each iteration $k$, we have $\delta_{k}=0$, i.e. $\nabla\Omega_{k}(\mathbf{x}_{k+1})=0$.
We relax 5.1 and discuss the inexact proximal steps in Section 7.
5.1 Convergence analysis
We discuss the convergence analysis here, and missing proofs can be found in Appendix A. The convergence analysis of Algorithm 1 revolves around the following quantities:
$\begin{array}[]{ccl}F_{k}&\coloneqq&{\mathbb{E}}\left[F(\mathbf{x}_{k})-F^{\star}\right]\,,\\ \Delta_{k}&\coloneqq&{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]\,,\\ R_{k}&\coloneqq&{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]\,.\end{array}$ (4)
$F_{k}$ quantifies the distance between the current objective value and the lower bound. $\Delta_{k}$ bounds the distance between the current gradient estimate and the full gradient. $R_{k}$ bounds the distance between two consecutive iterates. We start by giving a descent lemma on $\Delta_{k}$, which is key to analyzing the variance reduction effect of momentum. Similar statements can be found in [7] and [12].
Lemma 5.2.
Given 3.4, and for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$\Delta_{k+1}\leq(1-\gamma)\Delta_{k}+\frac{L^{2}}{\gamma}R_{k}+\gamma^{2}\sigma^{2}\,.$
The $(1-\gamma)$ factor in the above lemma is the key to show that $\Delta_{k}$ decreases in the optimization process. Next, we discuss the per-iteration descent of $F_{k}$.
Recall that in this section, we have 5.1, which is equivalent to the following stationarity assumption:
$\nabla\Omega_{k}(\mathbf{x}_{k+1})=\nabla\psi(\mathbf{x}_{k+1})+\mathbf{m}_{k}+M(\mathbf{x}_{k+1}-\mathbf{x}_{k})=0\,.$ (5)
When $\psi$ is convex and non-differentiable, $\nabla\Omega_{k}(\mathbf{x}_{k+1})$ is just a subgradient of $\Omega_{k}$ at $\mathbf{x}_{k+1}$ that equals to zero, whose existence is guaranteed as the optimality condition. In the following lemmas and theorems, we can simply replace the gradient of $\Omega$ and $\psi$ with such choices of the subgradients and obtain the same results for convex and non-differentiable $\psi$.
In Section 7, we give an approximate stationarity assumption, which is more realistic in practice and shows the same convergence guarantees.
Now we give the following descent lemma on $F_{k}$:
Lemma 5.3.
Given 3.1, 3.3, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$F_{k+1}\leq F_{k}-\frac{M-L}{4}R_{k}+\frac{\Delta_{k}}{M-L}\,.$
Remark 5.4.
The constants in Lemma 5.3 can be slightly improved if 3.1 is replaced by the convexity of $\psi$.
Proof.
By 3.3, we have:
$F(\mathbf{x}_{k+1})$ $=f(\mathbf{x}_{k+1})+\psi(\mathbf{x}_{k+1})$
$\leq f(\mathbf{x}_{k})+\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle+\frac{L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\psi(\mathbf{x}_{k+1})$
$=f(\mathbf{x}_{k})+\langle\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle+\psi(\mathbf{x}_{k+1})+\frac{M}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}$
$\quad-\frac{M-L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\langle\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle\,.$
Then we apply 3.1 and notice that $\Omega_{k}(\mathbf{x}_{k})=\psi(\mathbf{x}_{k})$:
$F(\mathbf{x}_{k+1})$ $\leq F(\mathbf{x}_{k})-\frac{M-L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\langle\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle\,.$
Now we apply Young’s equality and get
$F(\mathbf{x}_{k+1})$ $\leq F(\mathbf{x}_{k})-\frac{M-L}{4}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\frac{\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}}{M-L}\,.$
We get the desired result by subtracting $F^{\star}$ and taking expectation on both sides. ∎
Note that these will be enough if we aim to prove convergence in terms of the distance between the iterates, but not enough to guarantee a convergence in terms of ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k})\right\rVert^{2}\right]$. Therefore, we relate the distance between iterates and the norm of the gradient in the following lemma:
Lemma 5.5.
Given 3.3 and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$R_{k}\geq\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{3(M^{2}+L^{2})}-\frac{\Delta_{k}}{M^{2}+L^{2}}\,.$
Proof.
Note that we can split $\nabla F(\mathbf{x}_{k+1})$ in the following way:
$\nabla F(\mathbf{x}_{k+1})$ $=\nabla f(\mathbf{x}_{k+1})+\nabla\psi(\mathbf{x}_{k+1})$
$=\mathbf{m}_{k}+\nabla\psi(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\,.$
Then we have:
$\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}$ $\leq 3\left\lVert\mathbf{m}_{k}+\nabla\psi(\mathbf{x}_{k+1})\right\rVert^{2}+3\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+3\left\lVert\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\,.$
Now apply the stationarity condition (5) on the first term, and use 3.3 on the third term, we get:
$\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}$ $\leq 3(M^{2}+L^{2})\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+3\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\,.$
We get the desired result by rearranging and taking expectations on both sides. ∎
Now, we can piece together all of the above lemmas and consider the following Lyapunov function:
$\Phi_{k}\coloneqq F_{k}+a\Delta_{k}\,,$ (6)
where $a$ is a constant to be determined later. Note that $a$ does not have an impact on the algorithm itself, and it only shows up in the analysis. Now we give the following lemma on $\Phi_{k}$:
Lemma 5.6.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3L}{(M-L)},a\coloneqq\frac{3}{8L}$, and $M>4L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L}{4M^{2}}\sigma^{2}\,.$ (7)
Remark 5.7.
Note that the above lemma sets the momentum parameter $\gamma$ to be of the order $\Theta(\frac{L}{M})$.
We defer the proof to Appendix A. Now, with the above lemma and following the standard arguments, for a total number of $K$ iterations, and for $M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$, we have:
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}\left(\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\Phi_{0}}{K}\right)\,.$
This indicates that the final convergence guarantee depends on $\Phi_{0}$, i.e. the initialization, which we discuss in the next section.
5.2 Initialization and convergence guarantees
As mentioned in Remark 5.7, $a=\mathcal{O}(\frac{1}{L})$, therefore $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\Delta_{0}}{L})$. There are some subtle differences in how $\Delta_{0}$ can be initialized between the non-composite and composite cases.
Non-composite case:
When $\psi\equiv 0$, i.e. $F\equiv f$, we set the initial momentum $\mathbf{m}_{-1}\coloneqq-\frac{\gamma\mathbf{g}_{0}}{1-\gamma}$, then $\Delta_{0}={\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]=\left\lVert\nabla f(\mathbf{x}_{0})\right\rVert^{2}\leq\mathcal{O}(LF_{0})$, where the last inequality follows from 3.3 and the fact that $F\equiv f$. Therefore, in the non-composite case, we can initialize the parameters such that $\Phi_{0}=\mathcal{O}(F_{0})$.
Composite case:
The composite case is slightly trickier. We set $\mathbf{m}_{-1}\coloneqq\mathbf{g}_{0}$ and get that $\mathbf{m}_{0}=\mathbf{g}_{0}$. Hence $\Delta_{0}={\mathbb{E}}\left[\left\lVert\mathbf{g}_{0}-\nabla\psi(\mathbf{x}_{0})\right\rVert^{2}\right]\leq\sigma^{2}$. Therefore, $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\sigma^{2}}{L})$. When $\sigma^{2}=\mathcal{O}(LF_{0})$, we get the same $\Phi_{0}=\mathcal{O}(F_{0})$ as in the non-composite case.
Mini-batch initialization:
In the case that $\sigma^{2}$ is much larger than $F_{0}$, if we have access to a constant size (not depending on the target error) mini-batch initially, then we can set $\mathbf{g}_{0}=\frac{1}{b_{0}}\sum_{i=1}^{b_{0}}\nabla f(\mathbf{x}_{0},\xi_{i})$ where $b_{0}\coloneqq\frac{\sigma^{2}}{LF_{0}}$, i.e. $\mathbf{g}_{0}$ is a mini-batch stochastic gradient of batch size $\frac{\sigma^{2}}{LF_{0}}$. Then we have $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\sigma^{2}}{b_{0}L})=\mathcal{O}(F_{0})$, which is the same as in the non-composite case.
Now we state the final convergence guarantees of Algorithm 1:
Theorem 5.8.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3L}{(M-L)}$, there exists $M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ such that after at most
$K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$
iterations of Algorithm 1 it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\varepsilon$, where $\mathbf{x}_{\kappa}$ is chosen from $\kappa\in\left\{1,\cdots,K\right\}$ uniformly at random.
More specifically, if $\psi\equiv 0$, we can initialize $\mathbf{m}_{-1}$ such that $K=\mathcal{O}(\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon})$; Otherwise, we can initialize $\mathbf{m}_{-1}$ such that $K=\mathcal{O}(\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{\sigma^{4}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon}+\frac{\sigma^{2}}{\varepsilon})$; further, when an error-independent initial mini-batch is allowed, we can initialize $\mathbf{m}_{-1}$ with a $\frac{\sigma^{2}}{LF_{0}}$ batch such that $K=\mathcal{O}(\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon})$.
We note that the output criteria (selecting a random iterate) is standard in the literature and could also be replaced with an average of all iterates. In both cases, this can tracked efficiently (e.g. [33]).
When an initial mini-batch is not allowed, the convergence rate has a $\mathcal{O}(\frac{\sigma^{4}}{\varepsilon^{2}}+\frac{\sigma^{2}}{\varepsilon})$ term that is not present in the non-composite case. One natural question is whether this is an artifact of our analysis or inherent to the problem and the algorithm. Note that the first step of the Algorithm 1 coincides with the vanilla stochastic proximal gradient method. Consider the lower bound construction in Proposition 4.1: when starting at $\mathbf{x}_{0}=0$, which is a stationary point such that $F_{0}=0$, one such step would incur an $\mathcal{O}(\sigma^{2})$ error. Therefore, in the non-composite case without the initial mini-batch, the convergence rate must have some term that only depends on $\sigma^{2}$. In other words, the extra terms seem unavoidable.
6 Variance reduction effect of momentum
In Section 4, we demonstrated that the vanilla stochastic proximal gradient method cannot converge because the gradient noise variance keeps the iterates away from the stationary point of $F$. In this section, we show that the momentum term in Algorithm 1 can reduce the variance of the gradient noise at the same rate as the gradient norm. This is the key to the convergence of Algorithm 1 without batches. This variance reduction effect has been known in practice and implicitly used in the analysis of [7] and [12]. We precisely characterize such an effect in the composite optimization setting. Proofs are deferred to Appendix B.
We begin by refining the result of Lemma 5.6:
Lemma 6.1.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3\sqrt{2}L}{(M-L)},a\coloneqq\frac{\sqrt{2}}{8L}$, and $M>(1+3\sqrt{2})L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M}\Delta_{k}+\frac{27\sqrt{2}L}{4M^{2}}\sigma^{2}\,.$ (8)
With Lemma 6.1, we can now precisely quantify the variance reduction effect of the momentum method.
Theorem 6.2.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3\sqrt{2}L}{(M-L)}$, there exists $M=(1+3\sqrt{2})L+\frac{3\sqrt{3\sqrt{2}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ such that after at most
$K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$
iterations of Algorithm 1 it holds that ${\mathbb{E}}\left[\left\lVert\mathbf{m}_{\kappa}-\nabla f(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\varepsilon$, where $\kappa$ is chosen from $\kappa\in\left\{0,\cdots,K-1\right\}$ uniformly at random.
In words, the squared distance between the momentum $\mathbf{m}_{k}$ and the full gradient $\nabla f(\mathbf{x}_{k})$ decreases at the same rate as the squared norm of the gradient $\nabla F(\mathbf{x}_{k})$.
7 Inexact proximal step
While many existing works (e.g. [16, 14, 39, 17, 37, 42]) rely on the assumption that the proximal step can be solved exactly, this might not always be practically possible [3]. In this section, we briefly discuss extending our analysis to the case where the proximal step is solved inexactly and give an inexactness criterion similar to that of [41].
A crucial element in our previous analysis is the assumption that $\nabla\Omega_{k}(\mathbf{x}_{k+1})=0$, i.e. Equation 5. Therefore, defining the inexactness criteria as an approximate stationarity condition of $\Omega_{k}$ at $\mathbf{x}_{k+1}$ where $\delta_{k}\neq 0$ is natural. In particular, we define the criteria as follows:
${\mathbb{E}}\left[\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{M^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]}{16}+S_{k}\,,$ (9)
where $S_{k}$ will be decided later. Again, it should be understood that in the non-differentiable case $\nabla\Omega_{k}(\mathbf{x}_{k+1})$ is a certain subgradient of $\Omega_{k}$ at $\mathbf{x}_{k+1}$ such that Equation 9 holds.
Now we state the convergence result (proofs deferred to Appendix C):
Theorem 7.1.
Given 3.1, 3.3, and 3.4, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\sqrt{\frac{152}{17}}\frac{L}{M-L}$, and at iteration $k$,
${\mathbb{E}}\left[\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{M^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\right]}{16}+S_{k}\,,$
then there exists $M=4L+\sqrt{\frac{8KL\sigma^{2}}{\Phi_{0}}}$ such that after at most
$K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$
iterations of Algorithm 1 it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\frac{\varepsilon}{2}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$, where $\mathbf{x}_{\kappa}$ is chosen from $\kappa\in\left\{1,\cdots,K\right\}$ uniformly at random.
In particular, if for any $k$, $S_{k}\leq\frac{\varepsilon}{16}$, then ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\varepsilon$.
It remains to discuss how to minimize $\Omega_{k}$ such that the inexactness criteria (9) is satisfied. There is a line of research on minimizing gradient norm with SGD variants, e.g. [1], and one can also exploit the structure information in $\psi$ if there is any. Here, we state that SGD suffices for our purpose and give the following proposition, which is a simple modification of Proposition 2.6 in [41]:
Proposition 7.2.
If $\psi$ is convex and $\hat{L}$-smooth, and we have access to an unbiased gradient oracle of $\psi$ with variance at most $\hat{\sigma}^{2}$, then after at most
$T=\mathcal{O}\left((1+\frac{\hat{L}}{M})\ln{\frac{(\hat{L}+M)^{2}}{M^{2}}}+\frac{(\hat{L}+M)\hat{\sigma}^{2}}{MS_{k}}\right)$
iterations of SGD, the output $\hat{\mathbf{x}}$ of SGD satisfies the condition (9).
The proof of Proposition 7.2 is simple, and we refer interested readers to [41] for the discussions therein.
8 Experiments
In this section, we conduct numerical experiments to corroborate our theoretical findings and demonstrate the practical effectiveness of Algorithm 1.
8.1 Synthetic quadratic problem
Refer to caption Refer to caption Refer to caption
Figure 1: Comparison of Algorithm 1 and the vanilla stochastic proximal gradient method on the synthetic quadratic problem. The left, middle, and right figures correspond to $\sigma=5,25,125$, respectively. The vanilla stochastic proximal gradient method uses batch sizes $1,5,10$. The x-axis represents the number of gradient samples and is truncated to only show the first $10^{5}$ gradient samples.
We first consider a synthetic quadratic problem inspired by our lower bound construction in Proposition 4.1. We consider the following problem in $\mathbb{R}^{d}$: we set $f(\mathbf{x})=\frac{L\left\lVert\mathbf{x}\right\rVert^{2}}{2}$ and $\psi(\mathbf{x})=\frac{a\left\lVert\mathbf{x}\right\rVert^{2}}{2}$, where $a$ is sufficiently large. Note that the vanilla stochastic proximal gradient method and Algorithm 1 are oblivious to the parameter $a$. We add Gaussian noise to the gradients to control the stochastic level $\sigma^{2}$ of the stochastic oracle. We simulate the batch sample by dividing the variance by the batch size.
In Figure 1, we compare the performance of Algorithm 1 and the vanilla stochastic proximal gradient method. We set $d=5,L=1$ and $a=10^{4}$. We run the vanilla method with batch sizes $1,5,10$. We set $\sigma=5,25,125$ respectively. The parameter $M$ is tuned by a grid search in $\{1,10,100,10^{3},10^{4}\}$ for all methods, and the momentum parameter $\gamma$ is tuned by a grid search in $\{0.1,0.001,10^{-3},10^{-4},10^{-5}\}$. We set the maximum number of iterations to be $10^{4}$, and the tolerance is $0.02$. We see that as $\sigma^{2}$ increases, Algorithm 1 still reaches the desired tolerance, while the vanilla method with batch size $1$ fails to converge in all cases, and with batch sizes $5$ and $10$ only converges when $\sigma=5$. In particular, with batch size $1$, the error of vanilla method oscillates around $36$ with $\sigma^{2}=5^{2}$, around $590$ with $\sigma^{2}=25^{2}$, and around $9375$ with $\sigma^{2}=125^{2}$. In other words, the error of the vanilla method is indeed proportional to $\sigma^{2}$, as predicted by our lower bound result in Proposition 4.1. We also point out that the momentum method exhibits a jump in the error in the first several iterations, consistent with our analysis that the first step of the momentum method incurs an $\mathcal{O}(\sigma^{2})$ error as well.
8.2 Statistical preconditioning experiment
Now we consider the recent statistical preconditioner (proxy training) technique of [17, 41], where for some objective $\ell$, we consider $f=\ell-\hat{\ell}$ and $\psi=\hat{\ell}$. $\hat{\ell}$ is a “statistical preconditioner” defined on a sub-sample of the training dataset. Here we simulate the setup on Cifar-10 dataset [23]. $\ell$ is defined on the whole $50000$ training images, and $\hat{\ell}$ is defined on a subset of $2560$ training images. We follow the implementation of [41], with one difference: at each iteration $k$, [41] computes the full gradient ^$\nabla\hat{\ell}(\mathbf{x}_{k})$ while we only compute a stochastic gradient of batch size $128$. In the experiment, we use batch size of $512$ for $\ell$, and a batch size of $128$ for the SGD updates on $\Omega_{k}$. We perform a grid-search on the parameters $M$ and $\gamma$. The SGD on $\Omega_{k}$ takes $20$ iterations and a step-size $0.01$, which is tuned in [41]. The experiments demonstrate that SGD is effective and reliable for solving the proximal step with sufficient accuracy (see also the experiments in [41]). We see that the momentum method outperforms the vanilla method, and the convergence of the momentum method seems smoother than the vanilla method.
Figure 2: Comparison of Algorithm 1 and the vanilla stochastic proximal gradient method for the statistical preconditioning technique on Cifar-10 dataset. The left and right figures correspond to the training loss and test accuracy, respectively.
9 Conclusion
In this work we revisit the non-convex stochastic composite optimization problem, and address its convergence issue in the small batch regime. We show that the vanilla stochastic proximal method cannot converge to the stationary point beyond the variance of the gradient noise. We analyze the immensely successful Polyak momentum method in this context and establish its optimal convergence rate without any batch size requirement, demonstrating its superiority over the vanilla method. We conduct numerical experiments to corroborate our theoretical findings. In light of the past successes of proximal methods in ML, and the recent emerging application scenarios for proximal methods in DL our findings reinforce the robustness and the potential of the Polyak momentum method.
Appendix A Missing proofs in Section 5
We start by giving the proof of Lemma 5.2: See 5.2
Proof.
${\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $={\mathbb{E}}\left[\left\lVert(1-\gamma)\mathbf{m}_{k}+\gamma\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$={\mathbb{E}}\left[\left\lVert(1-\gamma)(\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k+1}))+\gamma(\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1}))\right\rVert^{2}\right]$
$=(1-\gamma)^{2}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+{\mathbb{E}}\left[\gamma^{2}\left\lVert\mathbf{g}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\leq(1-\gamma)^{2}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})+\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
Where the in the second equality we used the fact that $\mathbf{g}_{k+1}$ is unbiased. By Young’s inequality, for any $\alpha>0$, we have:
${\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]$ $\leq(1-\gamma)^{2}(1+\alpha){\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]$
$\quad+(1-\gamma)^{2}(1+\alpha^{-1}){\mathbb{E}}\left[\left\lVert\nabla f(\mathbf{x}_{k})-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
$\leq(1-\gamma)^{2}(1+\alpha){\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]$
$\quad+(1-\gamma)^{2}(1+\alpha^{-1})L^{2}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
where the last inequality follows from the smoothness of $f$. For $\alpha\coloneqq\frac{\gamma}{1-\gamma}$, we have:
${\mathbb{E}}\left[\left\lVert\mathbf{m}_{k+1}-\nabla f(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq(1-\gamma){\mathbb{E}}\left[\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\right]+\frac{L^{2}}{\gamma}{\mathbb{E}}\left[\left\lVert\mathbf{x}_{k}-\mathbf{x}_{k+1}\right\rVert^{2}\right]+\gamma^{2}\sigma^{2}$
Now we give the proof of Lemma 5.6:
See 5.6
Proof.
We put together Lemma 5.2 and Lemma 5.3 and get:
$\Phi_{k+1}$ $=F_{k+1}+a\Delta_{k+1}$
$\leq F_{k}-\frac{M-L}{4}R_{k}+\frac{\Delta_{k}}{M-L}$
$\quad+a\left[(1-\gamma)\Delta_{k}+\frac{L^{2}}{\gamma}R_{k}+\gamma^{2}\sigma^{2}\right]$
$=F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)R_{k}+a\gamma^{2}\sigma^{2}$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}$
If $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$, then we can plug Lemma 5.5 in and get:
$\Phi_{k+1}$ $\leq F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)(\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{3(M^{2}+L^{2})}-\frac{\Delta_{k}}{M^{2}+L^{2}})$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
$=F_{k}-\left(\frac{M-L}{12(M^{2}+L^{2})}-\frac{aL^{2}}{3\gamma(M^{2}+L^{2})}\right){\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}+\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
Now let
$\frac{M-L}{24(M^{2}+L^{2})}=\frac{aL^{2}}{3\gamma(M^{2}+L^{2})}$
then $a=\frac{\gamma(M-L)}{8L^{2}}$. Note that with this $a$, we have $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$. Now
$\frac{1}{a(M-L)}+\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}$ $=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{M-L}{8}\frac{1}{a(M^{2}+L^{2})}$
$=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{L^{2}}{\gamma(M^{2}+L^{2})}$
$=\frac{L^{2}}{\gamma}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)$
Therefore we need $\frac{L^{2}}{\gamma}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)\leq\gamma$, rearranging, it suffices to have:
$\gamma^{2}\geq L^{2}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)$
Note that $(M-L)^{2}\leq M^{2}+L^{2}$, it suffices to set $\gamma^{2}=\frac{9L^{2}}{(M-L)^{2}}$, or equivalently:
$\gamma=\frac{3L}{(M-L)}$
Note that this requires $M>4L$. Now
$a=\frac{M-L}{8L^{2}}\frac{3L}{M-L}=\frac{3}{8L}$
Putting them all together,
$\Phi_{k+1}\leq\Phi_{k}-\frac{M-L}{24(M^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L}{8(M-L)^{2}}\sigma^{2}$
Let $M=\tau L$ for some $\tau>4$, then
$\Phi_{k+1}\leq\Phi_{k}-\frac{\tau-1}{24(\tau^{2}+1)L}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27}{8(\tau-1)^{2}L}\sigma^{2}$
Note that for $\tau>4,\frac{\tau-1}{\tau^{2}+1}\geq\frac{1}{2\tau}$ and $\frac{1}{(\tau-1)^{2}}\leq\frac{2}{\tau^{2}}$. Therefore, for $M>(1+3\sqrt{3})L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L}{4M^{2}}\sigma^{2}$
Rearranging and summing Equation 7 from 00 to $K-1$, we get
$\frac{1}{48M}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\Phi_{0}+\frac{27KL\sigma^{2}}{4M^{2}}$
Therefore
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{48M\Phi_{0}}{K}+\frac{324L\sigma^{2}}{M}$
Recall that $M>4L$, then for $M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$. We have:
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}(\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\Phi_{0}}{K})$
Therefore, after at most $\mathcal{O}(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon})$ iterations, we have $\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\varepsilon$.
Appendix B Missing proofs in Section 6
See 6.1
Proof.
In the proof of Lemma 5.6, if we set
$\frac{L^{2}}{\gamma}\left(\frac{8}{(M-L)^{2}}+\frac{1}{M^{2}+L^{2}}\right)\leq\frac{\gamma}{2},$
or equivalently,
$\gamma=\frac{3\sqrt{2}L}{M-L},$
Note that this requires $M>(1+3\sqrt{2})L$. Now we have $a=\frac{\sqrt{2}}{8L}$ and:
$\Phi_{k+1}\leq\Phi_{k}-\frac{M-L}{24(M^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}L}{2(M-L)}\Delta_{k}+\frac{27\sqrt{2}L}{4(M-L)^{2}}\sigma^{2}$
Let $M=\tau L$ for some $\tau>(1+3\sqrt{2})$, then:
$\Phi_{k+1}\leq\Phi_{k}-\frac{\tau-1}{24(\tau^{2}+1)L}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2(\tau-1)}\Delta_{k}+\frac{27\sqrt{2}}{4(\tau-1)^{2}L}\sigma^{2}$
Note that for $\tau>(1+3\sqrt{2}),\frac{\tau-1}{\tau^{2}+1}\geq\frac{1}{2\tau},\frac{1}{2(\tau-1)}\geq\frac{1}{2\tau}$ and $\frac{1}{(\tau-1)^{2}}\leq\frac{2}{\tau^{2}}$. Therefore, for $M>(1+3\sqrt{2})L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\frac{3\sqrt{2}}{2M}\Delta_{k}+\frac{27\sqrt{2}L}{4M^{2}}\sigma^{2}$
Appendix C Missing proofs in Section 7
First, we notice that under 3.1 and 3.3, Lemma 5.3 still holds in the inexact case. Now we give an analogous result to Lemma 5.5 in the inexact case:
Lemma C.1.
Given 3.3, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$R_{k}\geq\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{4(M^{2}+L^{2})}-\frac{\Delta_{k}}{M^{2}+L^{2}}-\frac{{\mathbb{E}}\left[\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{M^{2}+L^{2}}$
Proof.
$\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}$ $=\left\lVert\nabla f(\mathbf{x}_{k+1})+\nabla\psi(\mathbf{x}_{k+1})\right\rVert^{2}$
$=\left\lVert\mathbf{m}_{k}+\nabla\psi(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\right\rVert^{2}$
$=\left\lVert M(\mathbf{x}_{k}-\mathbf{x}_{k+1})+\nabla\Omega(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\right\rVert^{2}$
$\leq 4M^{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+4\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}+4\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+4\left\lVert\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k})\right\rVert^{2}$
$\leq 4(M^{2}+L^{2})\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+4\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}+4\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}$
Rearranging, we have:
$-\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}\leq-\frac{\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}}{4(M^{2}+L^{2})}+\frac{\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}}{M^{2}+L^{2}}+\frac{\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}}{M^{2}+L^{2}}$
Notice that Lemma 5.2 still holds. We continue to use the same Lyapunov function $\Phi_{k}\coloneqq F_{k}+a\Delta_{k}$.
Lemma C.2.
Given 3.1, 3.3, and 3.4, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\sqrt{\frac{456}{17}}\frac{L}{M-L},a\coloneqq\sqrt{\frac{19}{408}}\frac{1}{L}$, that in each iteration we have:
$\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\leq\frac{M^{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}}{16}+S_{k},$
then if $M>7L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{68M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{12L}{M^{2}}\sigma^{2}+\frac{2}{17M}S_{k}$ (10)
Proof.
First we plug Equation 9 into Lemma C.1 and get:
$-R_{k}$ $\leq-\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{4(M^{2}+L^{2})}+\frac{\Delta_{k}}{M^{2}+L^{2}}+\frac{{\mathbb{E}}\left[\left\lVert\nabla\Omega(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{M^{2}+L^{2}}$
$\leq-\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{4(M^{2}+L^{2})}+\frac{\Delta_{k}}{M^{2}+L^{2}}+\frac{R_{k}}{16}+\frac{S_{k}}{M^{2}+L^{2}}$
Rearranging, we get:
$-R_{k}\leq-\frac{4{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{17(M^{2}+L^{2})}+\frac{16\Delta_{k}}{17(M^{2}+L^{2})}+\frac{16S_{k}}{17(M^{2}+L^{2})}$ (11)
Recall in the proof of Lemma 5.6, we have:
$\Phi_{k+1}$ $\leq F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)R_{k}+a\gamma^{2}\sigma^{2}$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}$
If $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$, then we can plug Equation 11 in and get:
$\Phi_{k+1}$ $\leq F_{k}-\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\left(\frac{4{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{17(M^{2}+L^{2})}-\frac{16\Delta_{k}}{17(M^{2}+L^{2})}-\frac{16S_{k}}{17(M^{2}+L^{2})}\right)$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
$=F_{k}-\frac{4}{17}\left(\frac{M-L}{4(M^{2}+L^{2})}-\frac{aL^{2}}{\gamma(M^{2}+L^{2})}\right){\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]$
$\quad+\left(1-\gamma+\frac{1}{a(M-L)}+\frac{16}{17}\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}\right)a\Delta_{k}+a\gamma^{2}\sigma^{2}$
$\quad+\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{16S_{k}}{17(M^{2}+L^{2})}$
Now let
$\frac{M-L}{8(M^{2}+L^{2})}=\frac{aL^{2}}{\gamma(M^{2}+L^{2})}$
then $a=\frac{\gamma(M-L)}{8L^{2}}$. Note that with this $a$, we have $\frac{M-L}{4}-\frac{aL^{2}}{\gamma}>0$. Now
$\frac{1}{a(M-L)}+\frac{16}{17}\left(\frac{M-L}{4}-\frac{aL^{2}}{\gamma}\right)\frac{1}{a(M^{2}+L^{2})}$ $=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{2(M-L)}{17}\frac{1}{a(M^{2}+L^{2})}$
$=\frac{8L^{2}}{\gamma(M-L)^{2}}+\frac{16L^{2}}{17\gamma(M^{2}+L^{2})}$
$=\frac{8L^{2}}{\gamma}\left(\frac{1}{(M-L)^{2}}+\frac{2}{17(M^{2}+L^{2})}\right)$
Therefore we need $\frac{8L^{2}}{\gamma}\left(\frac{1}{(M-L)^{2}}+\frac{2}{17(M^{2}+L^{2})}\right)\leq\gamma$, rearranging, it suffices to have:
$\gamma^{2}\geq 8L^{2}\left(\frac{1}{(M-L)^{2}}+\frac{2}{17(M^{2}+L^{2})}\right)$
Note that $(M-L)^{2}\leq M^{2}+L^{2}$, it suffices to set $\gamma^{2}=\frac{152}{17}\frac{L^{2}}{(M-L)^{2}}$, or equivalently:
$\gamma=\sqrt{\frac{152}{17}}\frac{L}{M-L}$
Note that this requires $M>\frac{(\sqrt{152}+\sqrt{17})L}{\sqrt{17}}$, where it suffices that $M>4L$. Now
$a=\frac{M-L}{8L^{2}}\sqrt{\frac{152}{17}}\frac{L}{M-L}=\sqrt{\frac{19}{156}}\frac{1}{L}$
Putting them all together,
$\Phi_{k+1}\leq\Phi_{k}-\frac{M-L}{34(M^{2}+L^{2})}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{4L}{(M-L)^{2}}\sigma^{2}+\frac{2(M-L)S_{k}}{17(M^{2}+L^{2})}$
Let $M=\tau L$ for some $\tau>4$, then
$\Phi_{k+1}\leq\Phi_{k}-\frac{\tau-1}{34(\tau^{2}+1)L}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{4}{(\tau-1)^{2}L}\sigma^{2}+\frac{2(\tau-1)}{17(\tau^{2}+1)L}S_{k}$
Note that for $\tau>4,\frac{\tau-1}{\tau^{2}+1}\geq\frac{1}{2\tau},\frac{1}{(\tau-1)^{2}}\leq\frac{2}{\tau^{2}}$ and $\frac{\tau-1}{\tau^{2}+1}\leq\frac{1}{\tau}$. Therefore, for $M>7L$:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{68M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{8L}{M^{2}}\sigma^{2}+\frac{2}{17M}S_{k}$
Rearranging and summing Equation 10 over $k=0,\ldots,K-1$, we get:
$\frac{1}{68M}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\Phi_{0}+\frac{8KL}{M^{2}}\sigma^{2}+\frac{2}{17M}\sum_{k=0}^{K-1}S_{k}$
Therefore,
$\frac{1}{K}\sum_{k=0}^{K=1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\frac{68M\Phi_{0}}{K}+\frac{544L\sigma^{2}}{M}+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$
Recall that $M>4L$, then for $M=4L+\sqrt{\frac{8KL\sigma^{2}}{\Phi_{0}}}$, we have:
$\frac{1}{K}\sum_{k=0}^{K=1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}\left(\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\sigma^{2}}{K}\right)+\frac{8}{K}\sum_{k=0}^{K-1}S_{k}$
Appendix D Proximal gradient mapping
In this section, we briefly comment on the proximal gradient mapping, which is the convergence criteria used in most existing works [16, 14, 39, 17, 37, 42]. Proximal gradient mapping is defined in a very general context, where we consider the constrained composite optimization problem:
$\min_{\mathbf{x}\in X}F(\mathbf{x})\coloneqq f(\mathbf{x})+\psi(\mathbf{x}),$
with the $1$-strongly convex mirror map $r:X\to\mathbb{R}$. For any vector $\mathbf{g}$ and $\mathbf{x}$, and scalar $M>0$, define:
$\mathbf{x}^{+}\coloneqq\operatorname*{arg\,min}_{\mathbf{x}^{\prime}\in X}\{\langle\mathbf{g},\mathbf{x}^{\prime}-\mathbf{x}\rangle+\psi(\mathbf{x})+\beta_{r}(\mathbf{x}^{\prime},\mathbf{x})\}$
where $\beta_{r}(\mathbf{x}^{\prime},\mathbf{x})=r(\mathbf{x}^{\prime})-r(\mathbf{x})-\langle\nabla r(\mathbf{x}),\mathbf{x}^{\prime}-\mathbf{x}\rangle$ is the Bregman divergence of $r$. Then the proximal gradient mapping is defined as:
$P_{X}(\mathbf{x},\mathbf{g},M)\coloneqq M(\mathbf{x}-\mathbf{x}^{+}),$
and in the literature, the convergence is studied in terms of $\left\lVert P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)\right\rVert^{2}$. We point out that it is very easy to prove an analogous version of Lemma 5.3 and Lemma 5.5 in terms of $\left\lVert P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)\right\rVert^{2}$ (which implies that our results extend to the proximal gradient mapping case), but we omit the details here. Instead, we argue that the size of $\nabla F(\mathbf{x})$ is a more natural convergence criterion. Consider the simplest situation where $\psi$ is convex and differentiable, $X=\mathbb{R}^{d}$, and the mirror map $r(\mathbf{x})=\frac{1}{2}\left\lVert\mathbf{x}\right\rVert^{2}$ (i.e. the Euclidean geometry). Then we have the usual stationarity condition for $\mathbf{x}^{+}$:
$\nabla f(\mathbf{x})+\nabla\psi(\mathbf{x}^{+})+M(\mathbf{x}^{+}-\mathbf{x})=0.$
Therefore, we have:
$P_{X}(\mathbf{x},\nabla f(\mathbf{x}),M)=\nabla f(\mathbf{x})+\nabla\psi(\mathbf{x}^{+}).$
In other words, the proximal gradient mapping gives the gradient of $f$ at the current point plus the gradient of $\psi$ at the next point. In contrast, in our analysis, we directly consider the gradient of $F$ at the next point, and we believe that this is more natural and intuitive.